Tim’s Bicycle Shop sells 21-speed bicycles. For purposes of a cost-volume-profit analysis, the shop owner has divided sales into two categories, as follows:

Product Type Sales Price Invoice Cost Sales Commission
High-quality $ 1,300 $ 700 $ 80
Medium-quality 750 420 60

Three-quarters of the shop’s sales are medium-quality bikes. The shop’s annual fixed expenses are $159,600. (In the following requirements, ignore income taxes.)

Required:

  1. Compute the unit contribution margin for each product type.

  2. What is the shop’s sales mix?

  3. Compute the weighted-average unit contribution margin, assuming a constant sales mix.

  4. What is the shop’s break-even sales volume in dollars? Assume a constant sales mix.

  5. How many bicycles of each type must be sold to earn a target net income of $99,750? Assume a constant sales mix.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Tim’s Bicycle Shop
Workings
Calculation of contribution margin per unit High-quality Medium-quality Note
Sales Price                1,300.00                               750.00 A
Less:
Invoice Cost                   700.00                               420.00
Sales Commission                      80.00                                 60.00
Contribution per unit                   520.00                               270.00 C
Calculation of sales mix High-quality Medium-quality Note
Balance of units sold 1/4 3/4
Sales mix 25% 75% D Because 1/4 is 5% and 3/4 is 75%.
Calculation of weighted-average unit contribution margin High-quality Medium-quality Total Note
Contribution per unit                   520.00                               270.00 See C
Sales mix 25% 75% See D
Weighted-average Contribution per unit                   130.00                               202.50               332.50 E=C*D
Calculation of break-even sales units Total Note
Annual fixed expenses            159,600.00 F
Weighted-average Contribution per unit                   332.50 See E
Break-even sales units                   480.00 G=F/E
See A See D See G I=D*G J=A*I
Calculation of shop’s break-even sales volume in dollars Sell Price Sales Mix % Total Break-even Units Break-even Units Sales Value
High-quality                1,300.00 25%               480.00                 120.00       156,000.00
Medium-quality                   750.00 75%               480.00                 360.00       270,000.00
Total       426,000.00
So shop’s break-even sales volume is $ 426,000.
If target net income $ 99,750:
Calculation of target even sales units Note
Annual fixed expenses            159,600.00
Target net income              99,750.00
Target Contribution           259,350.00 L
Weighted-average Contribution per unit                   332.50 See E
Target sales units                   780.00 M=L/E
See D See M N=D*M
Calculation of break-even sales units- Product wise Sales Mix % Total Break-even Units Break-even Units
High-quality 25%                               780.00               195.00
Medium-quality 75%                               780.00               585.00
So, 195 units of High-quality cycles and 585 units Medium-quality cycles should be sole to earn target net income of $ 99,750.